OpenStudy (anonymous):

Help please solve the equation 8X-2(X+1)=2(3X-1) 0 2 no solution identity

OpenStudy (debbieg):

Distribute through the ( )'s on each side. Then combine all like terms (x terms and non-x terms).

OpenStudy (anonymous):

what does the distribute through the ()s mean

OpenStudy (okra77):

identity all real numbers

OpenStudy (anonymous):

Thank you Okra77!

OpenStudy (debbieg):

Is means, distribute.... 8X-2(X+1)=8X-2X-2 A(B+C)=AB+AC